OpenStudy (anonymous):

how do you solve 6p+9=5p

OpenStudy (raden):

first, subtract by 5p to both sides 6p+9-5p = 5p - 5p simplify, will be : p+9 = 0

OpenStudy (raden):

then subtract both sides by -9 p+9-9 = 0-9 p = -9

OpenStudy (raden):

i mean by 9

OpenStudy (anonymous):